# Learning efficient erasure protocols for an underdamped memory

### v1

@@ Abstract
Appendix S1 Langevin equation non-dimensionalization and numerical integration
The first oscillation mode of the cantilever is modeled by an harmonic oscillator of mass $m$, stiffness $k$, and damping coefficient $\gamma$. It is subject to a feedback electrostatic force $F_{\rm e}$ and thermal noise due to the contact with the thermostat. The dimensionful position $\bar{x}$ of the cantilever is described by the Langevin equation
$m\frac{d^{2}\bar{x}}{d\bar{t}^{2}}=-k\bar{x}-\gamma\frac{d\bar{x}}{d\bar{t}}+F_{e}+\sqrt{2k_{B}T\gamma}\xi,$ (S4)
with $\bar{t}$ the dimensionful time. The electrostatic force can take only two values in our feedback implementation, which we set as ±$\pm k\bar{x}_{1}$ with a proper definition of the origin of $\bar{x}$. The natural scales of this problem are $\sigma=\sqrt{k_{\rm B}T/k}$ for length, $\omega_{0}^{-1}=\sqrt{m/k}$ for time, and $k_{\rm B}T$ for energy. We therefore introduce $\bar{x}=\sigma x$ and $\bar{t}=\omega_{0}^{-1}t$ into Eq. S4 in order to derive the dimensionless Eq. 2, with the quality factor of the simple harmonic oscillator being $Q=m\omega_{0}/\gamma$.
To numerically integrate Eq. 2 we introduce the cantilever velocity $v=\dot{x}$, giving
$\dot{v}+Q^{-1}v=-U^{\prime}(x)+\sqrt{2Q^{-1}}\,\xi.$ (S5)
Writing the left-hand side of (S5) as ${\rm e}^{-t/Q}\partial_{t}\left(v{\rm e}^{t/Q}\right)$, and assuming that $x$ is constant over the interval $[t,t+\Delta t]$, we multiply (S5) by ${\rm e}^{t/Q}$, integrate over the interval $[t,t+\Delta t]$ and divide both sides by ${\rm e}^{(t+\Delta t)/Q}$ to obtain
$v(t+\Delta t)-\alpha v(t)=-(1-\alpha)QU^{\prime}(x)+\eta(t),$ (S6)
@@
$\eta(t)\equiv{\rm e}^{-(t+\Delta t)/Q}\sqrt{2Q^{-1}}\int_{t}^{t+\Delta t}{\rm d}t^{\prime}\,\xi(t^{\prime}){\rm e}^{t^{\prime}/Q}.$ (S7)
Since $\xi(t)$ is normally distributed so is $\eta(t)$, with correlations $\langle\eta(t)\rangle=0$, $\langle\eta(t)\eta(t^{\prime})\rangle=0$, and $\langle\eta(t)^{2}\rangle=1-\alpha^{2}$. Eq. (S6) is the second line of Eq. 3, upon making the replacement (with minor abuse of notation) $\eta(t)\to\sqrt{1-\alpha^{2}}\xi(t)$. The correlations of $\xi$ are as given. The first line of Eq. 3 follows from the integration of $\dot{v}=x$ over $[t,t+\Delta t]$.
Appendix S2 Benchmarking the learning algorithm
Figure S5: Trap-translation problem of Ref. 22 for an underdamped particle. We show trap position as a function of time (top) and work distributions (bottom) resulting from three protocols: (a) linear; (b) linear with jumps; and (c) a protocol expressed by a neural network trained by genetic algorithm to minimize $\langle W\rangle$. The latter protocol is close to optimal. The black dashed line in the top panels is the optimal protocol (S9). The black dashed vertical line in the lower panels is the optimal mean work value $W^{\star}$, Eq. (S10), and the dashed distribution is that resulting from the optimal protocol. Here averages and distributions are calculated from $5\times 10^{5}$ independent trajectories. The initial velocity $v(0)$ and the initial position $x(0)$ are both chosen from a Gaussian distribution with zero mean and unit variance.
In this section we demonstrate the ability of the evolutionary learning procedure used in this work to identify efficient protocols for underdamped systems (we have previously demonstrated the same for overdamped passive- and active-matter systems [29, 31, 32]).
@@
$\lambda^{\star}(t)=\Lambda(t+Q^{-1})+\Lambda\left[\delta(t)-\delta(t-t_{\rm f})\right],$ (S9)
@@
$W^{\star}=\frac{\lambda_{\rm f}^{2}}{2+Qt_{\rm f}}.$ (S10)
@@
To simulate the optimal protocol (S9) we must modify Eq. 3: if we write $U^{\prime}(x)$ as $U_{0}^{\prime}(x)+\Lambda\left[\delta(t)-\delta(t-t_{\rm f})\right]$, where $U_{0}^{\prime}(x)$ is the force due to the first term in Eq. (S9), then the second line of Eq. 3 becomes
$v(t+\Delta t)={\rm r.h.s}+\Lambda\left(\delta_{t,0}-\delta_{t,t_{\rm f}}\right).$ (S11)
@@
Here “${\rm r.h.s}$” denotes the right-hand side of Eq. 3 (with $U^{\prime}(x)$ replaced by $U_{0}^{\prime}(x)$), and $\delta_{i,j}$ is the Kronecker delta, equal to unity if $i=j$ and equal to zero otherwise. The work associated with the delta kicks is $\Delta W=\Lambda(v_{\rm i}-v_{\rm f})$, where $v_{\rm i}$ is the velocity immediately before the first kick and $v_{\rm f}$ is the velocity immediately after the second kick.
@@
This problem poses a difficult task for any learning algorithm, because the optimal protocol contains delta functions. Moreover, the delta functions are important: in Fig. S5(a) we show the results of a linear protocol $\lambda(t)=\lambda_{\rm f}t/t_{\rm f}$, which produces mean work $\langle W\rangle\approx 3.71$, and in Fig. S5(b) we show the result of a linear protocol with jumps, given by Eq. (S9) without the delta-function terms, which produces mean work $\langle W\rangle\approx 3.25$. Both values of work are considerably in excess of the optimal value.
@@
In Fig. S5(c) we show the protocol expressed by a deep neural network trained by genetic algorithm to minimize $\langle W\rangle$, following the procedure described in Ref. 29. This protocol has jumps at the start and end, and rapidly-varying features near the start and end. It produces mean work $\langle W\rangle\approx 0.66$, only about 4% larger than the optimal value. The distribution of work produced by the neural-network protocol is similar to that of the optimal distribution.
@@
The evolutionary learning algorithm has therefore identified a regularized version of the delta-function impulses seen in the optimal protocol, resulting in a protocol associated with a mean work only slightly larger than optimal. This comparison establishes confidence in the ability of the learning algorithm to identify efficient protocols for underdamped systems. In this work we show that it can identify protocols for underdamped systems that are considerably more efficient than those designed by hand. The results of this section suggest that the rapidly-varying features seen in Fig. 3(b-c) may become delta-function impulses in the true optimal limit, but also that the differences between the learned protocols and the hypothetical optimal ones are likely to be small.
Appendix S3 Training details
@@
We train the neural-network protocols following the procedure used in Ref. 29, specified by the code deposited in Ref. 34. Briefly, we simulate $10^{4}$ independent trajectories of the simulation model using a protocol $[x_{0}(t),x_{1}(t)]=[0,5]+{\bm{g}}_{\bm{\theta}}(t/t_{\rm f})$, where ${\bm{g}}$ is the two-dimensional output of a deep neural network whose input is $t/t_{\rm f}$ and whose weights are ${\bm{\theta}}$. From these $10^{4}$ trajectories we construct an order parameter $\phi$ (described below). We then adjust the weights ${\bm{\theta}}$ using a genetic algorithm instructed to extremize $\phi$. To do so we run 50 lots of $10^{4}$ independent trajectories, each controlled by a neural network with slightly different randomly-chosen values of ${\bm{\theta}}$. We select the best 5 neural networks (the 5 associated with the best values of $\phi$), and clone and mutate these networks in order to produce a new set of 50 neural networks. Mutations consist of adding Gaussian random numbers $\epsilon\sim{\mathcal{N}}(0,10^{-4})$ to each weight of the neural network. We repeat this procedure until $\phi$ stops evolving, with $n_{\rm ev}$ used to indicate the number evolutionary steps taken.
@@
$\phi=\begin{cases}\langle(\Delta x)^{2}\rangle+50&\text{if }P_{\rm s}(1)<0.9;\\ 1-P_{\rm s}(1)+\langle E_{\rm tot}\rangle/100&\text{otherwise}.\end{cases}$ (S12)
Here $\langle(\Delta x)^{2}\rangle=\langle(x(t_{\rm f})+5)^{2}\rangle$ is the mean-squared distance between the final position of the cantilever and the final position ($-x_{1}(t_{\rm f})=-5$) of the center of the left-hand potential well, assuming erasure to state 0. For erasure to state 1, we use the same protocol with the sign of $x_{0}(t)$ reversed. $P_{\rm s}(1)$ is the mean success rate for a single erasure (erasure to state 0 is deemed successful if $x(t_{\rm f})<0$), and $\langle E_{\rm tot}\rangle$ is the mean total energy $E_{\rm tot}=\frac{1}{2}v(t_{\rm f})^{2}+U[x(t_{\rm f})]$ at the end of the protocol. The first clause of Eq. (S12) encourages the neural network to produce protocols in which the cantilever position ends close to the final position of the left-hand potential well. When the probability of erasure is greater than 90%, the second clause becomes active and encourages protocols that 1) perform erasure with high fidelity, and 2) leave the cantilever with small values of kinetic and potential energy.
Fig. S6 shows the evolution of $\langle(\Delta x)^{2}\rangle$ (top), $P_{\rm s}(1)$ (middle), and $\langle E_{\rm tot}\rangle$ (bottom) as a function of the number $n_{\rm ev}$ of evolutionary steps. For each value of $n_{\rm ev}$ we show the values of these order parameters associated with the neural network that gives rise to the smallest value of $\phi$. The number of trajectories required after $n_{\rm ev}$ steps is $50\times 10^{4}\times n_{\rm ev}$. The trained protocol is then used in experiment.
@@
Figure S6: Evolution of the components of the order parameter (S12) as a function of the number $n_{\rm ev}$ of evolutionary steps, during the training of the neural-network protocol shown in Fig. 1(c).
Appendix S4 Additional data
Figure S7: Probability distributions $\rho(x,t)$ in the format of Fig. 1(c) resulting from the four protocols considered in this work, in simulation (left) and experiment (right). Panel (a) shows the basic protocol of Fig. 1(c). Panel (b) shows the enhanced protocol of Fig. 3(a). Panel (c) shows the learned protocol of Fig. 3(b), designed to maximize erasure fidelity and minimize total energy upon performing a single erasure. Panel (d) shows the learned protocol of Fig. 3(c), designed to maximize the reliability of multiple successive erasures. In all cases we show a single erasure to state 0, starting from equilibrium [26]. Movies of the time evolution $\rho(x,t)$ for all protocols are available as Supplemental Material [27].
In Fig. S7 we show the probability distributions $\rho(x,t)$ in the format of Fig. 1(c) and Fig. 3 resulting from the four protocols. In Fig. S8 we show time-ordered snapshots in the format of Fig. 1(d) for three protocols. In all cases, simulation and experiment are consistent.
Figure S8: Time-ordered snapshots for the basic and enhanced protocols, and for the learned protocol of Fig. 3(c). We show the potential $U(x)$, the associated Boltzmann distributions $\rho_{0}(x)$, and the distribution of cantilever tip positions $\rho(x)$ in experiment (green) and simulation (black dashed) for a single erasure to state 0 [26]. The enhanced and learned protocols achieve erasure with respective probabilities of 99.2% and 99.9%; the main difference between these protocols lies in the distribution of kinetic energies after erasure, as shown in Fig. 4. Movies of the time evolution of $U(x,t)$ and $\rho(x,t)$ for all protocols are available as Supplemental Material [27].

### v2

@@ Acknowledgements.
Appendix A Langevin equation non-dimensionalization and numerical integration
The first oscillation mode of the cantilever is modeled by an harmonic oscillator of mass $m$, stiffness $k$, and damping coefficient $\gamma$. It is subject to a feedback electrostatic force $F_{\mathrm{el}}$ and thermal noise due to the contact with the thermostat. The dimensionful position $\bar{x}$ of the cantilever is described by the Langevin equation
$m\frac{d^{2}\bar{x}}{d\bar{t}^{2}}=-k\bar{x}-\gamma\frac{d\bar{x}}{d\bar{t}}+F_{\mathrm{el}}+\sqrt{2k_{B}T\gamma}\xi,$ (4)
with $\bar{t}$ the dimensionful time. The electrostatic force can take only two values in our feedback implementation, which we set as ±$\pm k\bar{x}_{1}$ with a proper definition of the origin of $\bar{x}$. The natural scales of this problem are $\sigma=\sqrt{k_{\rm B}T/k}$ for length, $\omega_{0}^{-1}=\sqrt{m/k}$ for time, and $k_{\rm B}T$ for energy. We therefore introduce $\bar{x}=\sigma x$ and $\bar{t}=\omega_{0}^{-1}t$ into Eq. 4 in order to derive the dimensionless Eq. (2), with the quality factor of the simple harmonic oscillator being $Q=m\omega_{0}/\gamma$.
To numerically integrate Eq. (2) we introduce the cantilever velocity $v=\dot{x}$, giving
$\dot{v}+Q^{-1}v=-U^{\prime}(x,t)+\sqrt{2Q^{-1}}\,\xi.$ (5)
Writing the left-hand side of (5) as ${\rm e}^{-t/Q}\partial_{t}\left(v{\rm e}^{t/Q}\right)$, and assuming that $x$ is constant over the interval $[t,t+\Delta t]$, we multiply (5) by ${\rm e}^{t/Q}$, integrate over the interval $[t,t+\Delta t]$ and divide both sides by ${\rm e}^{(t+\Delta t)/Q}$ to obtain
$v(t+\Delta t)-\alpha v(t)=-(1-\alpha)QU^{\prime}(x,t)+\eta(t),$ (6)
@@ Appendix A Langevin equation non-dimensionalization and numerical integration
$\eta(t)\equiv{\rm e}^{-(t+\Delta t)/Q}\sqrt{2Q^{-1}}\int_{t}^{t+\Delta t}{\rm d}t^{\prime}\,\xi(t^{\prime}){\rm e}^{t^{\prime}/Q}.$ (7)
Since $\xi(t)$ is normally distributed so is $\eta(t)$, with correlations $\langle\eta(t)\rangle=0$, $\langle\eta(t)\eta(t^{\prime})\rangle=0$, and $\langle\eta(t)^{2}\rangle=1-\alpha^{2}$. Eq. (6) is the second line of Eq. (3), with $\tilde{\xi}(t)=\eta(t)/\sqrt{1-\alpha^{2}}$ a random noise from a centered normal distribution of unit variance. The first line of Eq. (3) follows from the integration of $\dot{v}=x$ over $[t,t+\Delta t]$.
Appendix B Benchmarking the learning algorithm
Figure 5: Trap-translation problem of Ref. 22 for an underdamped particle. We show trap position as a function of time (top) and work distributions (bottom) resulting from three protocols: (a) linear; (b) linear with jumps; and (c) a protocol expressed by a neural network trained by genetic algorithm to minimize $\langle W\rangle$. The latter protocol is close to optimal. The black dashed line in the top panels is the optimal protocol (9). The black dashed vertical line in the lower panels is the optimal mean work value $W^{\star}$, Eq. (10), and the dashed distribution is that resulting from the optimal protocol. Here averages and distributions are calculated from $5\times 10^{5}$ independent trajectories. The initial velocity $v(0)$ and the initial position $x(0)$ are both chosen from a Gaussian distribution with zero mean and unit variance.
In this section we demonstrate the ability of the evolutionary learning procedure to identify efficient protocols for underdamped systems (we have previously demonstrated the same for overdamped passive- and active-matter systems [30, 32, 33]).
@@ Appendix B Benchmarking the learning algorithm
$\lambda^{\star}(t)=\Lambda(t+Q^{-1})+\Lambda\left[\delta(t)-\delta(t-t_{\rm f})\right],$ (9)
@@
$W^{\star}=\frac{\lambda_{\rm f}^{2}}{2+Qt_{\rm f}}.$ (10)
@@
To simulate the optimal protocol (9) we must modify Eq. (3): if we write $U^{\prime}(x,t)$ as $U_{0}^{\prime}(x,t)+\Lambda\left[\delta(t)-\delta(t-t_{\rm f})\right]$, where $U_{0}^{\prime}(x,t)$ is the force due to the first term in Eq. (9), then the second line of Eq. (3) becomes
$v(t+\Delta t)={\rm r.h.s}+\Lambda\left(\delta_{t,0}-\delta_{t,t_{\rm f}}\right).$ (11)
@@
Here “${\rm r.h.s}$” denotes the right-hand side of Eq. (3) (with $U^{\prime}(x,t)$ replaced by $U_{0}^{\prime}(x,t)$), and $\delta_{i,j}$ is the Kronecker delta, equal to unity if $i=j$ and equal to zero otherwise. The work associated with the delta kicks is $\Delta W=\Lambda(v_{\rm i}-v_{\rm f})$, where $v_{\rm i}$ is the velocity immediately before the first kick and $v_{\rm f}$ is the velocity immediately after the second kick.
@@
This problem poses a difficult task for any learning algorithm, because the optimal protocol contains delta functions. Moreover, the delta functions are important: in Fig. 5(a) we show the results of a linear protocol $\lambda(t)=\lambda_{\rm f}t/t_{\rm f}$, which produces mean work $\langle W\rangle\approx 3.71$, and in Fig. 5(b) we show the result of a linear protocol with jumps, given by Eq. (9) without the delta-function terms, which produces mean work $\langle W\rangle\approx 3.25$. Both values of work are considerably in excess of the optimal value.
@@
In Fig. 5(c) we show the protocol expressed by a deep neural network trained by genetic algorithm to minimize $\langle W\rangle$, following the procedure described in Ref. 30. This protocol has jumps at the start and end, and rapidly-varying features near the start and end. It produces mean work $\langle W\rangle\approx 0.66$, only about 4% larger than the optimal value. The distribution of work produced by the neural-network protocol is similar to that of the optimal distribution.
@@
The evolutionary learning algorithm has therefore identified a regularized version of the delta-function impulses seen in the optimal protocol, resulting in a protocol associated with a mean work only slightly larger than optimal. This comparison establishes confidence in the ability of the learning algorithm to identify efficient protocols for underdamped systems. In the main text we show that it can identify protocols for underdamped systems that are considerably more efficient than those designed by hand. The results of this section suggest that the rapidly-varying features seen in Fig. 3(b-c) may become delta-function impulses in the true optimal limit, but also that the differences between the learned protocols and the hypothetical optimal ones are likely to be small.
Appendix C Training details
@@
We train the neural-network protocols following the procedure used in Ref. 30, specified by the code deposited in Ref. 37. Briefly, we simulate $10^{4}$ independent trajectories of the simulation model using a protocol $[x_{0}(t),x_{1}(t)]=[0,5]+{\bm{g}}_{\bm{\theta}}(t/t_{\rm f})$, where ${\bm{g}}$ is the two-dimensional output of a deep neural network whose input is $t/t_{\rm f}$ and whose weights are ${\bm{\theta}}$. From these $10^{4}$ trajectories we construct an order parameter $\phi$ (described below). We then adjust the weights ${\bm{\theta}}$ using a genetic algorithm instructed to extremize $\phi$. To do so we run 50 lots of $10^{4}$ independent trajectories, each controlled by a neural network with slightly different randomly-chosen values of ${\bm{\theta}}$. We select the best 5 neural networks (the 5 associated with the best values of $\phi$), and clone and mutate these networks in order to produce a new set of 50 neural networks. Mutations consist of adding Gaussian random numbers $\epsilon\sim{\cal N}(0,10^{-4})$ to each weight of the neural network. We repeat this procedure until $\phi$ stops evolving, with $n_{\rm ev}$ used to indicate the number evolutionary steps taken.
Figure 6: Evolution of the components of the order parameter in Eq. (12) as a function of the number $n_{\rm ev}$ of evolutionary steps, during the training of the neural-network protocol shown in Fig. 1(c).
@@
$\phi=\begin{cases}\langle(\Delta x)^{2}\rangle+50&\text{if }P_{\rm s}(1)<0.9;\\ 1-P_{\rm s}(1)+\langle E_{\rm tot}\rangle/100&\text{otherwise}.\end{cases}$ (12)
Here $\langle(\Delta x)^{2}\rangle=\langle(x(t_{\rm f})+5)^{2}\rangle$ is the mean-squared distance between the final position of the cantilever and the final position ($-x_{1}(t_{\rm f})=-5$) of the center of the left-hand potential well, assuming erasure to state 0. For erasure to state 1, we use the same protocol with the sign of $x_{0}(t)$ reversed. $P_{\rm s}(1)$ is the mean success rate for a single erasure (erasure to state 0 is deemed successful if $x(t_{\rm f})<0$), and $\langle E_{\rm tot}\rangle$ is the mean total energy $E_{\rm tot}=\frac{1}{2}v(t_{\rm f})^{2}+U(x(t_{\rm f}),t_{f})$ at the end of the protocol. The first clause of Eq. (12) encourages the neural network to produce protocols in which the cantilever position ends close to the final position of the left-hand potential well. When the probability of erasure is greater than 90%, the second clause becomes active and encourages protocols that 1) perform erasure with high fidelity, and 2) leave the cantilever with small values of total energy.
Fig. 6 shows the evolution of $\langle(\Delta x)^{2}\rangle$ (top), $P_{\rm s}(1)$ (middle), and $\langle E_{\rm tot}\rangle$ (bottom) as a function of the number $n_{\rm ev}$ of evolutionary steps. For each value of $n_{\rm ev}$ we show the values of these order parameters associated with the neural network that gives rise to the smallest value of $\phi$. The number of trajectories required after $n_{\rm ev}$ steps is $50\times 10^{4}\times n_{\rm ev}$. The trained protocol is then used in experiment.
@@
Appendix D Additional data
Figure 7: Probability distributions $\rho(x,t)$ in the format of Fig. 1(c) resulting from the four protocols considered in this work, in simulation (left) and experiment (right). Panel (a) shows the basic protocol of Fig. 1(c). Panel (b) shows the enhanced protocol of Fig. 3(a). Panel (c) shows the learned protocol of Fig. 3(b), designed to maximize erasure fidelity and minimize total energy upon performing a single erasure. Panel (d) shows the learned protocol of Fig. 3(c), designed to maximize the reliability of multiple successive erasures. In all cases we show a single erasure to state 0, starting from equilibrium [25]. Movies of the time evolution $\rho(x,t)$ for all protocols are available as ancillary files [26].
In Fig. 7 we show the probability distributions $\rho(x,t)$ in the format of Fig. 1(c) and Fig. 3 resulting from the four protocols. In Fig. 8 we show time-ordered snapshots in the format of Fig. 1(d) for three protocols. In all cases, simulation and experiment are consistent.
Figure 8: Time-ordered snapshots for the basic and enhanced protocols, and for the learned protocol of Fig. 3(c). We show the potential $U(x,t)$, the associated Boltzmann distributions $\rho_{0}(x)$, and the distribution of cantilever tip positions $\rho(x,t)$ in experiment (green) and simulation (black dashed) for a single erasure to state 0 [25]. The enhanced and learned protocols achieve erasure with respective probabilities of 99.2% and 99.9%; the main difference between these protocols lies in the distribution of kinetic energies after erasure, as shown in Fig. 4. Movies of the time evolution of $U(x,t)$ and $\rho(x,t)$ for all protocols are available as ancillary files [26].